OpenStudy (king):

xy+x+y=-13 yz+y+z=-9 xz+x+z=5 Find x,y,z.

OpenStudy (anonymous):

x(y+1) + y = -13 so x = (-13 -y)/(y+1)

OpenStudy (anonymous):

substitute that in for x, and the rest is just a simple 2 equations with 2 unknown variable problem

OpenStudy (king):

please wait.....lemme try...

OpenStudy (anonymous):

try.

OpenStudy (king):

what is -13-y*z+1?

OpenStudy (anonymous):

can you be more clear? post your work

OpenStudy (king):

wait i got (-13z-13-yz-y)/y+1 +z=5

OpenStudy (anonymous):

ok use equation 2 now: y(z+1) + z = -9 so y = (-9-z)/(z+1) plug that in for y

OpenStudy (king):

i got z^2+2z=3

OpenStudy (anonymous):

ok. that is a quadratic equation. solve it

OpenStudy (king):

can z hav 2 values?

OpenStudy (anonymous):

yes

OpenStudy (king):

then ans is 1,-3

OpenStudy (king):

for z

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

now substitute for z in the other equations

OpenStudy (anonymous):

(for simplicity's sake, substitute into the original equations)

OpenStudy (king):

ok ,y =-5,3

OpenStudy (anonymous):

correct

OpenStudy (king):

x=-4,2

OpenStudy (anonymous):

there you go. That's it!

OpenStudy (king):

thnx....ure awesome!!!!!!!!!!!!!